OpenStudy (anonymous):

Block 1, of mass = 0.700 , is connected over an ideal (massless and frictionless) pulley to block 2, of mass , as shown. Assume that the blocks accelerate as shown with an acceleration of magnitude = 0.350 and that the coefficient of kinetic friction between block 2 and the plane is = 0.350. Find the mass of block 2, , when = 30.0.

OpenStudy (anonymous):

use this formula!!! for block of mass m1, there is two forces tension and the weight force... f=ma or t-w=m1a or t-m1g=m1a you will be able to find the tension in the string... as the string is connected the acceleration and the tension is same for both the block for block of mass m2, t-f(friction)=m2a then at last m2=tension /(acceleration+coeff.of kinetic friction *gravity) you will find the answer that is 1.879 kg notice that this answer is valid if u r taking S.I unit , you haven't mention any unit in your question...